OpenStudy (anonymous):

expand: (x+y^2)(x+c-y^2)

OpenStudy (anonymous):

multiply all the letters on the left bracket by each letter in the right bracket

OpenStudy (anonymous):

We have: \[ (x+y^2)(x+c-y^2)=(x+y^2)(x-y^2+c)=x^2-y^4+c(x+y^2)=\\ x^2-y^4+cx+cy^2 \]And, we are done.